Two equal resistors of 15 kΩ each are connected in series across a 120 V source. A voltmeter when connected across each read 48 V in each case. Determine the resistance of the voltmeter.


1
Expert's answer
2022-06-03T16:22:02-0400
"R_{eq}=\frac{15*R}{15+R}+15"

"I=\frac{V}{R_{eq}}=\frac{120}{\frac{15*R}{15+R}+15}"

"V_{ab}=120-I*15"

"48=120-\frac{120}{\frac{15*R}{15+R}+15}=120(1-\frac{15+R}{30R+225})"

"\frac{15+R}{30R+225}=0.6"

"18R-135=15+R"

"R=8.8\:\Omega"